OpenStudy (anonymous):

w=5 2 w+2 (w+5) = 30 plz explain

OpenStudy (anonymous):

hey whats up jessie :) basically when you substitute the 5 in for w, you will get 30. 2(5) + 2((5)+5) = 10 + 2(10) right? then finally, 10 +20 = 30 cheers

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

no probs jessie :)